What is the pH of a 0.473 M aqueous solution of hydrofluoric acid? Ka (HF) = 7.1x10-4
Answer #1

ŕ . Ka= 7018104 . HF Ht t 0.473nno 0.473- e initial final o ke ze- [ht] (P7) = 7-14104 ka= 0:473-22 = [HE] a z= 0018 = [ht] =